OpenStudy (anonymous):

when tossing 4 coins simultaneously , find the probability that at least 1 head is showing .

OpenStudy (mathmate):

Are you familiar with probability trees?

OpenStudy (anonymous):

no not really ! @mathmate

OpenStudy (mathmate):

Do you have any suggestions how to solve the problem?

OpenStudy (nurali):

The probability of getting tails for a given coin toss is (1/2). So for four coin tosses it is (1/2)^4. Thus: P(all 4 tails) = (1/2)*(1/2)*(1/2)*(1/2) = (1/2)^4 = 1/16 Hence: P(atleast one head) = 1 - P(all tails) = 1 - 1/16 = 15/16

OpenStudy (mathmate):

@anna113 @nurali has solved it exactly how I would have solved the problem, athough I may not have explained it the same way. I suggest you follow up with nurali if you need further explanations.

OpenStudy (anonymous):

thanks @mathmate and @Nurali

OpenStudy (nurali):

Anytime.

OpenStudy (mathmate):

You're welcome! Don't forget to give a medal to nurali! :)